Sol Summary – November 22nd

Sol Summary Report

November 22, 2017

A quiet day is a good day on Mars

Shannon Rupert, MDRS Director

Mission Status: Good.  Crew 183 has arrived and will begin their
one-week mission tomorrow.

Sol Activity Summary: We spent the morning and much of the afternoon
cleaning the Hab.  Late in the afternoon, we both took a break and did
a little exploration. Crew 183 arrived at the station at 7:30 pm.

Look Ahead Plan: Training tomorrow for Crew 183.

Anomalies in work:  Out of sim.

Weather:  Calm, sunny and wonderfully warm.

Crew Physical Status:  We are both fine.  I personally am very tired tonight.

EVA: Brandon went exploring to the south this afternoon, and I went
north to the southernmost branch of Tank Wash.  It was amazing.

Reports to be file: Sol Summary and Operations Report.

Support Requested: Will someone be standing by to add the crew to the
Mission Support eList?  I’m worried because it is Thanksgiving.

Operations Report – November 22nd

Crew 182+ Operations Report  22NOV2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Lower deck wall heater is now operational, but
forced air furnace continues to alternate between warm and cool air.

I keep forgetting to report than the porthole came out of the main
airlock again.

Notes on non-nominal systems: nothing to report

Generator (hours run): Ran until 9 am, then off until 5:45 pm

Solar— SOC 97 % when generator was turned on

Diesel –  not checked %

Propane –  not checked psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –   0 gallons.

Water (trailer) –  0 gallons.

Water (static) –  350 gallons

Trailer to Static Pump used –  no

Water (loft) – Static to Loft Pump used –  not known

Water Meter:  not checked

Toilet tank emptied: Yesterday

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) 350 s moved to their normal place

Oil Added? no

ATV Fuel Used:  not known Gals

# Hours the ATVs were Used today: 5 min

Notes on ATVs: The 300’s battery is low, did not jump today.

Deimos rover used:  yes


Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR, not yet ready


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used:  Not yet ready


Beginning charge:

Ending charge:

Currently charging: no

Opportunity rover used:  not yet ready


Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used:  no yet ready


Beginning charge:

Ending charge:

Currently charging: no

HabCar used? Yes, hauled 100 gallons of water and emptied the 100
gallons Brandon got yesterday

General notes and comments: Cleaning day

Summary of internet: not known

Summary of suits:  I checked all of them today.  Two more chargers not
functional.  Unfortunately one of the helmets was cracked in such a
way that it will not be repairable.  Removed from service.  One of the
backpacks is missing one of the two straps that hold the collar on the

Summary of Hab operations: Brandon and I cleaned the Hab.  It took all
morning and then some.  Looking good now.

Summary of GreenHab operations:  Watered.  Used 4 gallons.  Removed
hose from frost free faucet.  Crews should get water from faucet
without hose from now until spring to prevent freeze up.  I’ve been
thinking I may get the 55 gallon drum from town to keep water in
inside the GreenHab.  It might also increase humidity, which seems to
stay around 16 %.

Summary of ScienceDome operations: Nothing to report.

Summary of RAMM operations:  Nothing to report.

Summary of health and safety issues:  We are safe and we are healthy.

Questions, concerns and requests to Mission Support:  Nothing, thank you.

Copyright © The Mars Society. All rights reserved. | Main Site